# Disentangling Hippocampal Shape Variations: A Study of Neurological Disorders Using Mesh Variational Autoencoder with Contrastive Learning

### v2

@@ 2.3 Supervised Contrastive VAE (Ours)
We introduce a contrastive loss based on Frosst et al. (2019) applied to the excitation-inhibition mechanism inspired by Ding et al. (2020) and Kiechle et al. (2023) including a threshold hyperparameter in the regression loss for disentangling latent space of a VAE. The loss function $L_{contr}$ is composed of three parts: $L_{vae}$, $L_{contr}^{cls}$, and $L_{contr}^{reg}$ shown in 8. The first part is the loss function of the VAE, while the other two are the contrastive loss functions. Contrastive loss learns the representations of the data in the latent space, and our first contrastive loss function $L_{contr}^{cls}$ enforces the similarity between the samples of the same class and the dissimilarity between the samples of different classes. We use it to disentangle the first latent variable ($z_{1}$) that correlates with the bump (present or absent) in the torus dataset and disease (healthy or MS) in the hippocampus dataset. Here, the binary labels are represented by $y$. The loss is demonstrated in Figure 1.
@@
We formulate another similar loss function (equation 10) to address our regression problem of disentangling $z_{2}$ (depicted in figure 1) based on continuous labels such as ages or scales. We employ an extra threshold ($Th$) hyperparameter to transform the regression problem into a classification problem. The loss function categorizes data objects into the same class if their labels fall within a specified range ($|y_{i}-y_{j}|\leq Th$) determined by the threshold.
@@ 3.4.2 Ablation Study
Additionally, we conducted an ablation study to justify the selection of a mesh-based convolutional neural network.The plot in figure 4 demonstrates that varying the spiral sequence length affects the SAP score for both the hippocampus and torus datasets. The non-linear relationship between spiral length and SAP scores shows that optimal performance occurs at specific lengths, implying that neighborhood connectivity impacts disentanglement. This justifies our use of SpiralNet++, which leverages such connectivity effectively.
@@ 3.4.5 MS vs Hippocampal Volume Across Age Groups
Figure 7: Volume differences between healthy and MS hippocampus are shown using six age ranges. Volume differences for the left and right hippocampus are shown separately for all three supervised disentanglement methods (SG VAE, Attribute VAE, and our SC VAE). The plot depicts a higher right hippocampus volume difference across ages derived by all three methods. However, our method shows the highest volume difference, which is closer to the average trend found in (Valdés Cabrera et al., 2023). The volumes are calculated by generating 10 shapes (using the three models) from each range, and then taking the average.
@@ 4.1 Discussion
Despite the promising results, there are limitations to our approach. The generalization of our model to diverse populations and datasets needs further exploration. We also need to improve the disentanglement performance of disease status for better prediction and reconstruction. Additionally, the mesh convolution technique we used, as outlined by Gong et al. (2019), necessitates the registration of meshes to a template mesh. Consequently, it is crucial to investigate methods that do not rely on assumptions about mesh topology for the analysis of complex shapes. Future work would involve the incorporation of longitudinal data and exploration of the generalizability of the proposed method.

### v3

@@ 2.3 Supervised Contrastive VAE (Ours)
We introduce a contrastive loss based on Frosst et al. (2019) applied to the excitation-inhibition mechanism inspired by Ding et al. (2020) and Kiechle et al. (2023) including a threshold hyperparameter in the regression loss for disentangling latent space of a VAE. The loss function $L_{contr}$ is composed of three parts: $L_{vae}$, $L_{contr}^{cls}$, and $L_{contr}^{reg}$ shown in 8. The first part is the loss function of the VAE, while the other two are the contrastive loss functions.
Contrastive loss learns the representations of the data in the latent space, and our first contrastive loss function $L_{contr}^{cls}$ enforces the similarity between the samples of the same class and the dissimilarity between the samples of different classes. In our problem setting, classification inherently deals with discrete labels, where the model’s objective is to distinguish between distinct classes. The loss function $L_{contr}^{cls}$ encourages the latent space to separate data points based on these discrete labels. This loss only applies to the latent variable responsible for the classification task. We use it to disentangle the first latent variable ($z_{1}$) that correlates with the bump (present or absent) in the torus dataset and disease (healthy or MS) in the hippocampus dataset. Here, the binary labels are represented by $y$. The loss is demonstrated in Figure 1.
@@
On the other hand, Regression deals with continuous targets in our setting. We employ a threshold parameter (Th) in equation 10 to simulate a classification problem using continuous targets. This is a hyperparameter that controls the granularity of the regression task. It divides the values into separate bins and is completely different from the classification loss function. This loss only applies to the latent variable responsible for the regression task. We formulate the loss function to address our regression problem of disentangling $z_{2}$ (depicted in figure 1) based on continuous labels such as ages or scales. The loss function categorizes data objects into the same class if their labels fall within a specified range ($|y_{i}-y_{j}|\leq Th$) determined by the threshold.
@@ 3.4.2 Ablation Study
Additionally, we conducted an ablation study to verify the effect of neighborhood information on disentanglement in mesh-based convolutional neural networks. The plot in figure 4 demonstrates that varying the spiral sequence length affects the SAP score for both the hippocampus and torus datasets. The non-linear relationship between spiral length and SAP scores shows that optimal performance occurs at specific lengths, implying that neighborhood connectivity impacts disentanglement. Therefore, the length of the spiral, as neighborhood information, has a significant effect on disentanglement. On a scale of 1, the SAP score varies in the range of 0.07 and 0.06 for the torus and hippocampus data, respectively. The scores would not change much or remain similar if a reasonable spiral length does not affect disentanglement. This justifies our use of SpiralNet++, which leverages such connectivity effectively.
@@ 3.4.5 MS vs Hippocampal Volume Across Age Groups
Figure 7: Volume differences between healthy and MS hippocampus are shown using six age ranges. Volume differences for the left and right hippocampus are shown separately for all three supervised disentanglement methods (SG VAE, Attribute VAE, and our SC VAE). The plot depicts a higher right hippocampus volume difference across ages derived by all three methods. The volumes are calculated by generating 10 shapes (using the three models) from each range, and then taking the average. Our method shows the highest volume difference ($\sim$4.5%) among the methods and is closer (compared to other methods) to the average trend (9%) found in a study by Valdés Cabrera et al. (2023) that calculated hippocampus volumes from DTI.
@@ 4.1 Discussion
Despite the promising results, there are limitations to our approach. The generalization of our model to diverse populations and datasets needs further exploration. We also need to improve the disentanglement performance of disease status for better prediction and reconstruction. Additionally, the mesh convolution technique we used, as outlined by Gong et al. (2019), necessitates the registration of meshes to a template mesh. Consequently, it is crucial to investigate methods that do not rely on assumptions about mesh topology for the analysis of complex shapes. Also, Our classification and regression losses are mostly similar, with only minor differences. Therefore, a unified formulation could be utilized to compactly represent the loss function. Future work would involve the incorporation of longitudinal data and exploration of the generalizability of the proposed method.